If variable costs for producing 1 unit are 4.5 EUR, that for the production we need to lease a machinery for 3000 EUR per year, if the production exceeds 1000 units we have to pay workers additional 3000 EUR.

We have one demand for 1500 units for a price of 10 EUR per unit. Should we launch the product? how should we do the calculation?
Answer #1

Total cost = Fixed cost + Variable cost

Total revenue = 1500 * 10 = 15000

Lease payment for machine = 3000

Additional payment to labor = 3000

Total cost for producing 1500 units = variable cost * units + Lease payment for machine + Additional payment to labor

=  1.5*1500 + 3000 + 3000

= 2250 + 3000 + 3000 = 8250

Profit = TR - TC = 15000 - 8250 = 6750

As there is positive profit, product should be launched